OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (9, 7) and (4, –8). y −?? =?? (x − 9)

OpenStudy (anonymous):

@iGreen @Nnesha

OpenStudy (igreen):

Plug the points in the slope formula and find the slope.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-2 and -12

OpenStudy (igreen):

No..

OpenStudy (igreen):

-8 = y2 7 = y1 4 = x2 9 =x1

OpenStudy (anonymous):

and then divide right

OpenStudy (igreen):

Yep

OpenStudy (anonymous):

2 and 2

OpenStudy (igreen):

No.. \(\dfrac{-8 - 7}{4 - 9}\)

OpenStudy (anonymous):

-15/-5=3

OpenStudy (anonymous):

still confused DX

OpenStudy (igreen):

Correct.

OpenStudy (igreen):

So the slope is 3. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y - y_1 = 3(x - 9)\) So do you know what 'y1' is going to be?

OpenStudy (anonymous):

is it kinda the same thing when we figured out the other one.

OpenStudy (igreen):

?

OpenStudy (igreen):

Sort of

OpenStudy (igreen):

What is in the 'x1' place?

OpenStudy (anonymous):

9

OpenStudy (igreen):

Yes, so the point must have a '9' for the 'x'

OpenStudy (igreen):

Which one of our points has that?

OpenStudy (anonymous):

(9,7)

OpenStudy (igreen):

Correct, and what's the y-value?

OpenStudy (anonymous):

I gtg D: so cya

OpenStudy (igreen):

Oh, well the y-value is 7..so that goes at 'y1'. \(y - 7 = 3(x - 9)\)